# Advancing Grounded Multimodal Named Entity Recognition via LLM-Based Reformulation and Box-Based Segmentation

### v1

@@ V-A1 Datasets
Detailed statistics of the datasets used by different modules are shown in Table III. For the MNER task, we conduct experiments on two classic MNER datasets, i.e., Twitter-2015 and Twitter-2017. Twitter-2015 is constructed by Zhang et al. [2], and Twitter-2017 is constructed by Yu et al. [4] based on the SNAP dataset [3]. For the GMNER task, the experiments are based on the currently only Twitter-GMNER dataset [17]. and used for training and validating the VE and VG modules are subsets of the Twitter-GMNER dataset. For the SMNER task, the statistical results are based on our proposed Twitter-SMNER dataset (detailed in § III). Additionally, the base training sets of all datasets are constructed solely based on gpt-3.5-turbo. And 4 different versions of LLMs (i.e., vicuna-7b-v1.5 [56], vicuna-13b-v1.5 [56], llama-2-7b-chat-hf [57], llama-2-13b-chat-hf [57]) are used to implement data augmentation of the training data. All dev and test sets do not involve data augmentation and are constructed based on gpt-3.5-turbo.
@@ V-A3 Implementation Details
Without being specified, all training is based on a single 4090 GPU. The best models selected on their respective dev sets are used for evaluation on the test sets. Aligned with previous research [17, 40, 41], precision (Pre.), recall (Rec.), and F1 score are utilized for evaluating the model performance.
@@
For OFA, the learning rate is 3e-5, the batch size is 32 and the training epoch is 10. For SeqTR [24], as the pre-training weights of SeqTR are not available, we use the checkpoint after pre-training and 5 epochs of fine-tuning on RefCOCO [23] as initial weights. 777https://github.com/seanzhuh/SeqTR The maximum length of input text is 30, the learning rate is 5e-4, the batch size is 64. And visual encoder is YOLOv3 [63]. The training epoch is 10. The remaining settings are consistent with the VG fine-tuning strategy of OFA [27] and SeqTR [24].
@@ V-B1 Results on MNER
Table IV shows the comparison of the proposed MNER module and different baseline methods. For fair comparison, the same set of methods uses the same text encoder. All results are averages of three runs using different random seeds. We present experimental results for different LLMs as well as Mix versions. The datasets used for experiments with individual LLM versions are constructed by the respective LLM. The training set for the Mix version includes data from all 5 LLMs, while the dev and test sets are only constructed by ChatGPT. The results show that LLMs with smaller parameters generally provide lower quality knowledge than LLMs with larger parameters. Additional external knowledge in different styles for the same samples contributes to further enhancing the capabilities of the model. Obviously, the performance of our MNER module surpasses the two SoTA methods MoRe [15] and ITA [14]. This indicates that the quality of external knowledge retrieved from LLMs is higher than that retrieved from Wikipedia. We also observe that MoRe performs worse than simple baseline methods in some cases. This is because auxiliary knowledge derived from Wikipedia may be redundant or contain irrelevant content. The above experiments demonstrate the effectiveness of the proposed MNER module.
@@ V-B2 Results on GMNER
Table V shows the comparison between RiVEG and existing GMNER methods. Following Yu et al. [17], we also count the MNER and Entity Extraction & Grounding (EEG) results. EEG focuses solely on extracting entity-region pairs without considering entity types. RiVEG achieves significant advantages over baseline methods. Without any data augmentation involved, RiVEG outperforms the previous SoTA method H-index by 9.47%, 4.78%, and 7.61% on all three tasks, respectively. After data augmentation, the lead further extends to 10.65%, 6.21%, and 8.83%. This demonstrates the rationality and effectiveness of the proposed framework. We maximize the performance of MNER to mitigate the error propagation in GMNER task, leverage the characteristics of the VG method to naturally bypass region feature pre-extraction, and address the weak correlation between images and text by leveraging a dedicated VE module. Note that the modular design of RiVEG allows it to benefit from any subsequent independent module research and achieve further performance improvements.
@@ V-B3 Results on SMNER
Table V also shows the SMNER results of the proposed RiVEG framework. As there are no existing SMNER baselines, we extend three of the most advanced GMNER baseline methods in the same manner, leveraging the localization capability of GMNER methods to prompt SAM in form of boxes. We also count the Entity Extraction & Segmentation (EES) subtask, which only focuses on the correctness of extracted entity-mask pairs. The experimental results demonstrate the feasibility of combining the visual object localization capability of any GMNER method with the zero-shot image segmentation capability of SAM to accomplish the SMNER task. The performance loss from boxes to masks is minimal, with the F1 score of all methods decreasing by only about 2%-3%. Additionally, due to the stronger capabilities in MNER and GMNER of RiVEG, our approach demonstrates significant superiority in the SMNER task. Without involving any data augmentation, RiVEG outperforms the best baseline method H-index by 6.46% and 8.32% on EES and SMNER task. After data augmentation, the advantage further expands to 7.57% and 9.39%, respectively.
@@ V-C1 Ablation Study of MNER
In ablation experiments shown in Table VI, we remove manually annotated samples and the MSEA module used to guide LLMs respectively. The backbone of this experiment is XLM-RoBERTa-CRF, and the LLM is gpt-3.5-turbo. w/o Manually Annotated Samples means letting the LLM directly provide explanations without any in-context example prompts. w/o MSEA means using a fixed set of randomly selected manually annotated examples for prompting. First, the results indicate that the performance can be improved regardless of how the external knowledge is constructed. Secondly, compared with knowledge generated without guidance, the quality of content generated with guidance is higher, and further marginal gains can be achieved by using more relevant examples. But more in-context examples are not always better, as too many in-context examples may introduce noise.
@@ V-C2 Exploration of More Visual Variants of RiVEG
Table VII shows the performance of more RiVEG visual variants. We additionally select the classic multimodal pre-training method ALBEF [28] for the VE module and the classic VG method SeqTR [24] for the VG module. The experimental results indicate that even without considering any data augmentation, the weakest variant exhibits highly competitive results compared with the baseline method. After data augmentation, all visual variants outperform the baseline method. The combination of stronger methods typically yields better performance. And we also observe a significant performance gap between SeqTR and OFA, which may be attributed to the text representation method adopted by SeqTR. The GRU vocabulary of SeqTR is constructed solely based on the training and validation data in or , which may result in difficulties in handling out-of-vocabulary words during its inference process.
@@ V-C3 Exploration of More Text Variants of RiVEG
Table VIII explores additional text variants of RiVEG. We fix the VE and VG modules, then replace XLM-RoBERTa with BERT in the MNER module, and remove the external knowledge generated by LLMs. The main conclusions are as follows: 1) The weaker text encoder clearly results in weaker MNER performance, leading to more pronounced error propagation and a decrease in GMNER and SMNER performance. 2) The MNER results demonstrate that the introduction of auxiliary refined knowledge significantly enhances performance. This once again validates the effectiveness of the proposed MNER module. 3) Experiments with the BERT in the w/o Auxiliary Refined Knowledge group suggest that even when the MNER performance is nearly identical to that of the baseline, RiVEG still significantly outperforms the baseline in GMNER and SMNER. This underscores the effectiveness of reformulating the overall task. 4) The tests in the w/o Auxiliary Refined Knowledge group also validate the effectiveness of data augmentation. When facing similar MNER results, the VE and VG methods enhanced with data augmentation achieve better GMNER and SMNER results. 5) All 14 variants in Table VII and Table VIII further demonstrate that combining the localization capability of VG methods and the zero-shot image segmentation capability of SAM can effectively accomplish the SMNER task, with only minimal performance loss.
@@ V-C5 Named Entity Referring Expressions Analysis
Table IX shows the performance of VE and VG methods when faced with different types of the text input. The text input of and constructed from ChatGPT is split into named entities along with entity types, and entity expansion expressions. The results indicate that different methods are suitable for different types of text inputs. Unified frameworks like OFA [27] are less affected by fine-grained real-world named entities due to their extensive pre-training on a large amount of data. And the impact is more significant for task-specific methods like SeqTR [24]. Different methods exhibit different characteristics, but the proposed hybrid named entity referring expressions, which balance between coarse-grained and fine-grained representations, yield the best performance across all methods.
@@ V-C6 Comparison Between Text Query-Based and Box Prompt-Based Segmentation Masks Prediction
Table X shows the comparison between two different methods of obtaining segmentation masks. Specifically, the training data for the two text-query based methods consist of named entity referring expression-mask pairs. These methods directly output segmentation masks based on the input text queries. And the training data for the box prompt-based method consists of named entity referring expression-box pairs. This method first predicts the box corresponding to the text query through visual grounding, and then prompts SAM to obtain the segmentation mask of the main visual object within the box. The results show that the two-stage box prompt-based method outperforms the single-stage text query-based method significantly. This may be attributed to the insufficient training data in the SMNER dataset, which limits the performance of text query-based segmentation methods. VG methods have lower requirements on the amount of data, and the emergence of SAM makes it possible to obtain the segmentation mask of the main visual object from boxes. Therefore, at the current stage, the idea of obtaining segmentation masks based on box prompts may be more suitable for the SMNER task.
@@ V-C7 Sensitivity Analysis of IoU Threshold
We also evaluate the performance of four RiVEG variants in Table V under different IoU thresholds for the GMNER and SMNER tasks. Fig. 8 illustrates that as the IoU threshold increases, the performance of all variants gradually decreases. The performance of GMNER is less affected by the increase in IoU threshold compared with SMNER. Additionally, the comparison between RiVEG and RiVEG further demonstrates the effectiveness of data augmentation. After data augmentation, the performance of RiVEG is nearly identical to that of RiVEG without data augmentation at the 0.5 threshold, indicating that data augmentation enhances the overall capability of the framework. Furthermore, RiVEG with data augmentation is more adapted to higher IoU thresholds, indicating that data augmentation endows the VG module with more accurate visual localization capability.
@@ V-C8 Case Study
VI Conclusion

### v2

@@ V-A1 Datasets
Detailed statistics of the datasets used by different modules are shown in Table IV. For the MNER task, we conduct experiments on two classic MNER datasets, i.e., Twitter-2015 and Twitter-2017. Twitter-2015 is constructed by Zhang et al. [2], and Twitter-2017 is constructed by Yu et al. [7] based on the SNAP dataset [6]. For the GMNER task, the experiments are based on the currently only Twitter-GMNER dataset [21]. and used for training and validating the VE and VG modules are subsets of the Twitter-GMNER dataset. For the SMNER task, the statistical results are based on our proposed Twitter-SMNER dataset (detailed in § III). Additionally, the base training sets of all datasets are constructed solely based on gpt-3.5-turbo. And 4 different versions of LLMs (i.e., vicuna-7b-v1.5 [59], vicuna-13b-v1.5 [59], llama-2-7b-chat-hf [60], llama-2-13b-chat-hf [60]) are used to implement data augmentation of the training data. These models are widely adopted in related studies [61, 62, 63] and are commonly regarded as representative choices. They span different scales (7B and 13B) and can be run on a single 24GB GPU, striking a balance between performance and reproducibility. All dev and test sets do not involve data augmentation and are constructed based on gpt-3.5-turbo.
@@ V-A3 Implementation Details
Without being specified, all training is based on a single 4090 GPU. The best models selected on their respective dev sets are used for evaluation on the test sets. Aligned with previous research [44, 45], precision (Pre.), recall (Rec.), and F1 score are utilized for evaluating the model performance.
@@
For OFA, the learning rate is 3e-5, the batch size is 32 and the training epoch is 10. For SeqTR [29], as the pre-training weights of SeqTR are not available, we use the checkpoint after pre-training and 5 epochs of fine-tuning on RefCOCO [28] as initial weights. 666https://github.com/seanzhuh/SeqTR The maximum length of input text is 30, the learning rate is 5e-4, the batch size is 64 and the training epoch is 10. The remaining settings are consistent with the VG fine-tuning strategy of OFA [32] and SeqTR [29].
@@ V-B1 Results on MNER
Table V shows the comparison of the proposed MNER module and different baseline methods. For fair comparison, the same set of methods uses the same text encoder. All results are averages of three runs using different random seeds. We present experimental results for different LLMs as well as Mix versions. The datasets used for experiments with individual LLM versions are constructed by the respective LLM. The training set for the Mix version includes data from all 5 LLMs, while the dev and test sets are only constructed by ChatGPT. The results show that LLMs with smaller parameters generally provide lower quality knowledge than LLMs with larger parameters. Additional external knowledge in different styles for the same samples contributes to further enhancing the capabilities of the model. Obviously, the performance of our MNER module surpasses the two SoTA methods MoRe [17] and ITA [16]. This indicates that the quality of external knowledge retrieved from LLMs is higher than that retrieved from Wikipedia. We also observe that MoRe performs worse than simple baseline methods in some cases. This is because auxiliary knowledge derived from Wikipedia may be redundant or contain irrelevant content. The above experiments demonstrate the effectiveness of the proposed MNER module.
@@ V-B2 Results on GMNER
Table VI shows the comparison between RiVEG and existing GMNER methods. Following Yu et al. [21], we also count the MNER and Entity Extraction & Grounding (EEG) results. EEG focuses solely on extracting entity-region pairs without considering entity types. RiVEG achieves significant advantages over baseline methods. Without any data augmentation involved, RiVEG outperforms the previous SoTA method H-index by 9.47%, 4.78%, and 7.61% on all three tasks, respectively. After data augmentation, the lead further extends to 10.65%, 6.21%, and 8.83%. This demonstrates the rationality and effectiveness of the proposed framework. We maximize the performance of MNER to mitigate the error propagation in GMNER task, leverage the characteristics of the VG method to naturally bypass region feature pre-extraction, and address the weak correlation between images and text by leveraging a dedicated VE module. Note that the modular design of RiVEG allows it to benefit from any subsequent independent module research and achieve further performance improvements.
@@ V-B3 Results on SMNER
Table VI also shows the SMNER results of the proposed RiVEG framework. As there are no existing SMNER baselines, we extend three of the most advanced GMNER baseline methods in the same manner, leveraging the localization capability of GMNER methods to prompt SAM in form of boxes. We also count the Entity Extraction & Segmentation (EES) subtask, which only focuses on the correctness of extracted entity-mask pairs. The experimental results demonstrate the feasibility of combining the visual object localization capability of any GMNER method with the zero-shot image segmentation capability of SAM to accomplish the SMNER task. The performance loss from boxes to masks is minimal, with the F1 score of all methods decreasing by only about 2%-3%. Additionally, due to the stronger capabilities in MNER and GMNER of RiVEG, our approach demonstrates significant superiority in the SMNER task. Without involving any data augmentation, RiVEG outperforms the best baseline method H-index by 6.46% and 8.32% on EES and SMNER task. After data augmentation, the advantage further expands to 7.57% and 9.39%, respectively.
@@ V-C1 Ablation Study of MNER
In ablation experiments shown in Table VII, we remove manually annotated samples and the MSEA module used to guide LLMs respectively. The backbone of this experiment is XLM-RoBERTa-CRF, and the LLM is gpt-3.5-turbo. w/o Manually Annotated Samples means letting the LLM directly provide explanations without any in-context example prompts. w/o MSEA means using a fixed set of randomly selected manually annotated examples for prompting. First, the results indicate that the performance can be improved regardless of how the external knowledge is constructed. Secondly, compared with knowledge generated without guidance, the quality of content generated with guidance is higher, and further marginal gains can be achieved by using more relevant examples. But more in-context examples are not always better, as too many in-context examples may introduce noise.
@@ V-C2 Exploration of More Visual Variants of RiVEG
Table VIII shows the performance of more RiVEG visual variants. We additionally select the classic multimodal pre-training method ALBEF [33] for the VE module and the classic VG method SeqTR [29] for the VG module. The experimental results indicate that even without considering any data augmentation, the weakest variant exhibits highly competitive results compared with the baseline method. After data augmentation, all visual variants outperform the baseline method. The combination of stronger methods typically yields better performance. And we also observe a significant performance gap between SeqTR and OFA, which may be attributed to the text representation method adopted by SeqTR. The GRU vocabulary of SeqTR is constructed solely based on the training and validation data in or , which may result in difficulties in handling out-of-vocabulary words during its inference process.
@@ V-C3 Exploration of More Text Variants of RiVEG
Table IX explores additional text variants of RiVEG. We fix the VE and VG modules, then replace XLM-RoBERTa with BERT in the MNER module, and remove the external knowledge generated by LLMs. The main conclusions are as follows: 1) The weaker text encoder clearly results in weaker MNER performance, leading to more pronounced error propagation and a decrease in GMNER and SMNER performance. 2) The MNER results demonstrate that the introduction of auxiliary refined knowledge significantly enhances performance. This once again validates the effectiveness of the proposed MNER module. 3) Experiments with the BERT in the w/o Auxiliary Refined Knowledge group suggest that even when the MNER performance is nearly identical to that of the baseline, RiVEG still significantly outperforms the baseline in GMNER and SMNER. This underscores the effectiveness of reformulating the overall task. 4) The tests in the w/o Auxiliary Refined Knowledge group also validate the effectiveness of data augmentation. When facing similar MNER results, the VE and VG methods enhanced with data augmentation achieve better GMNER and SMNER results. 5) All 14 variants in Table VIII and Table IX further demonstrate that combining the localization capability of VG methods and the zero-shot image segmentation capability of SAM can effectively accomplish the SMNER task, with only minimal performance loss.
@@ V-C5 Named Entity Referring Expressions Analysis
Table X shows the performance of VE and VG methods when faced with different types of the text input. The text input of and constructed from ChatGPT is split into named entities along with entity types, and entity expansion expressions. The results indicate that different methods are suitable for different types of text inputs. Unified frameworks like OFA [32] are less affected by fine-grained real-world named entities due to their extensive pre-training on a large amount of data. And the impact is more significant for task-specific methods like SeqTR [29]. Different methods exhibit different characteristics, but the proposed hybrid named entity referring expressions, which balance between coarse-grained and fine-grained representations, yield the best performance across all methods.
@@ V-C6 Comparison Between Text Query-Based and Box Prompt-Based Segmentation Masks Prediction
Table XI shows the comparison between two different methods of obtaining segmentation masks. Specifically, the training data for the two text-query based methods consist of named entity referring expression-mask pairs. These methods directly output segmentation masks based on the input text queries. And the training data for the box prompt-based method consists of named entity referring expression-box pairs. This method first predicts the box corresponding to the text query through visual grounding, and then prompts SAM to obtain the segmentation mask of the main visual object within the box. The results show that the two-stage box prompt-based method outperforms the single-stage text query-based method significantly. This may be attributed to the insufficient training data in the SMNER dataset, which limits the performance of text query-based segmentation methods. VG methods have lower requirements on the amount of data, and the emergence of SAM makes it possible to obtain the segmentation mask of the main visual object from boxes. Therefore, at the current stage, the idea of obtaining segmentation masks based on box prompts may be more suitable for the SMNER task.
@@ V-C7 Sensitivity Analysis of IoU Threshold
We also evaluate the performance of four RiVEG variants in Table VI under different IoU thresholds for the GMNER and SMNER tasks. Fig. 9 illustrates that as the IoU threshold increases, the performance of all variants gradually decreases. The performance of GMNER is less affected by the increase in IoU threshold compared with SMNER. Additionally, the comparison between RiVEG and RiVEG further demonstrates the effectiveness of data augmentation. After data augmentation, the performance of RiVEG is nearly identical to that of RiVEG without data augmentation at the 0.5 threshold, indicating that data augmentation enhances the overall capability of the framework. Furthermore, RiVEG with data augmentation is more adapted to higher IoU thresholds, indicating that data augmentation endows the VG module with more accurate visual localization capability.
@@ V-C8 Case Study
Despite the overall effectiveness of the box prompt-based segmentation approach, we observe a few rare failure cases where the predicted segmentation mask deviates from the expected visual object within the bounding box. As shown in Fig. 12, the model erroneously captures background regions or unrelated visual content. These examples reveal the limitations of box prompt-based segmentation under complex visual conditions. Although such errors are infrequent and do not significantly impact the overall performance, they highlight promising directions for future work, such as incorporating more adaptive or confidence-aware segmentation strategies.
VI Research Outlook
As illustrated in Fig. 13, related research suggests two potentially promising paradigms for the SMNER task:
1) The end-to-end paradigm leverages multimodal large language models to directly process multimodal inputs and yield structured outputs comprising the entity, its type, and corresponding segmentation. This paradigm offers potential advantages in unified optimization and task generalization, and notably, it is inherently free from error propagation between sequential modules. However, it typically requires substantial training resources and large-scale annotated data. Potential future directions for this paradigm include effective data augmentation strategies, optimized model architectures, and the exploration of using visual entity signals to enhance textual entity recognition in a reverse manner.
2) The cascading paradigm decomposes the task into multiple stages: first, the MNER module is applied to recognize textual entities, followed by diverse visual segmentation modules to localize entity masks. This design enables better modularity and interpretability, although it often introduces architectural complexity and potential error propagation across modules. Future directions include improving module capabilities and simplifying the architecture by reducing component count to mitigate error propagation.
This work adopts the cascading design, which demonstrates controllability and explainability. In the future, with the continued advancement of multimodal foundation models, task-specific end-to-end frameworks tailored for fine-grained entity-level understanding may become increasingly feasible, offering a compelling direction for further exploration.
VII Conclusion